OpenStudy (anonymous):

Solve the following equation, explaining what you did in each step of your solution. After solving the equation, show the steps to check your solution. 12 + 5(3 + 8x) = 267

OpenStudy (anonymous):

12 + 15 + 40x = 267

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

40x = 240

OpenStudy (anonymous):

i need the steps

OpenStudy (anonymous):

27 + 40x = 267...combine like terms

OpenStudy (anonymous):

move terms to other side making it 40x = 267 - 27

OpenStudy (anonymous):

40x = 240

OpenStudy (anonymous):

divide both sides by 40 to get x

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np